# Utilizing 5G NR SSB Blocks for Passive Detection and Localization of Low-Altitude Drones

### v2

@@ III-C Target detection
The detection decision is based on threshold, $\gamma$. The detector is given as $p\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\gamma$, where, the detector indicates the presence of a target when $p>\gamma$ and the absence of a target if $p<\gamma$. Here, $p$ is defined as $p=\max_{r}\text{PAF}$.
@@ IV Cramér-Rao Bound
For simplicity, we consider the CRB of the random variables, range and velocity, denoted as $\Theta$$=$$[\Theta_{i},\Theta_{j}]^{T}$$=$$[d,v]^{T}$, in a single target case ($K$$=$$1$). We evaluate this for a given signal-to-noise ratio at BsB, $\text{SNR}_{r}=\frac{\rho\beta g^{2}}{\sigma_{n}^{2}M}$, assuming a given RCS and fixed flight path. The received signal at the $n$th row, $l$th column of (7), disregarding the subscript $r,1$ and $k$, can be rewritten as
@@
where $g$$=$$\mathbf{a}^{T}(\theta_{d},\phi_{d})\mathbf{a}^{*}(\theta_{r},\phi_{r})$ is the beamforming gain of the $r$th transmitting beam with respect to the target’s angle and is bounded by $M(\text{dB})$--$3.7\text{dB}<\left\lvert g\right\rvert\leq M(\text{dB})$ irrespective of whether the target is located inside the grid of orthogonal beams, given that the $r$th beam is directed closest to the target among all $R$ beams. Formulate a vector $\mathbf{y}$$=$$[y_{1,1},y_{2,1},\dots,y_{2,1},y_{2,2},\dots,y_{N,L}]^{T},\;\mathbf{y}\in\mathbb{C}^{NL}$, from (9), i.e., by flattening $\mathbf{Y}$, where $[\mathbf{Y}]_{n,l}$$=$$y_{n,l}$. The Fisher information, $\text{FIM}(\Theta)$, for a complex Gaussian signal, as given in [11, Eq. (15.52)], is defined as
@@
where the first term in [11, Eq. (15.52)] is zero, $\mathbf{C_{y}}$$=$$\sigma_{n}^{2}\mathbf{I}$ is the covariance matrix of $\mathbf{y}$, since $\alpha$ remains constant within an OFDM block length $L$ in the Swerling 1 model, and $\mathbf{\mu_{y}}$$=$$\sqrt{\frac{\rho}{M}\beta}\alpha g[e^{-j2\pi f_{\Delta}\frac{d}{c}}e^{j4\pi\frac{v}{\lambda_{c}}T_{s}},\dots,e^{-j2\pi Nf_{\Delta}\frac{d}{c}}e^{j4\pi\frac{v}{\lambda_{c}}LT_{s}}]^{T}$ is the mean of $\mathbf{y}$.
@@ V Simulation Results
In Fig. 3, a target is generated with a random range, $d\leq d_{u}$, and random velocity, $\left\lvert v\right\rvert\leq v_{u}$. Fig. 3a shows CRB and range root mean square error (RMSE) of an $N^{\prime}$-point IFFT at SNRr$=$$-10$ dB, where the SSB line represents the received signal nullified at some symbols in the block, as shown in Fig. 1(a). Similarly, Fig. 3b shows the CRB and velocity RMSE of $L^{\prime}$-point FFT at SNRr$=$$-7$ dB. The results show that RMSE values for $N$$=$$240$, $L$$=$$4$, and the SSB block are approximately the same for both cases, range and velocity. The absence of some symbols in the SSB block has minimal effect compared to the completed block.
@@
Furthermore, the detection performance is investigated. The accuracy, $P_{c}$, is defined as $P_{c}$$=$$\frac{TP+TN}{T_{a}}$, where $TP$ is the number of true positives, $TN$ is the number of true negatives, and $T_{a}$ is the total number of experiments. We also incorporate a power-saving scheme by randomly deactivating a percentage of surveillance beams uniformly. A $100$% beam deactivation indicates that all surveillance beams are turned off, while a 00% deactivation means all beams remain active. Fig. 4 shows that with appropriate thresholds, $\gamma$$=$$4$ and $\gamma$$=$$6$, $P_{c}$ is maintained for up to $20$% beam deactivation. Note that the probabilities of the presence and absence of a target are both $0.5$.
Figure 4: $P_{c}$ of beam deactivation at different $\gamma$.
@@ VI Conclusions
This paper investigates target sensing using the SSB with a grid of orthogonal beams. Detection and localization are based on the 2D-FFT, assisted by PAF. The SSB symbols corresponding to the surveillance beams are used as an illuminator in a bistatic setup, with the base station as the sensing node. The simulation results for the given scenario show potential for UAV-like target sensing.
@@ Appendix A Proof of proposition 1
where $[\text{FIM}(\Theta)]_{v,d}$$=$$[\text{FIM}(\Theta)]_{d,v}$, $\mathbb{E}\{\left\lvert\alpha\right\rvert^{2}\}$$=$$1$, $\text{SNR}_{r}$$=$$\frac{\rho\beta g^{2}}{\sigma_{n}^{2}M}$, $\sum_{n=1}^{N}n$$=$$\frac{N(N+1)}{2}$ ,$\sum_{n=1}^{N}n^{2}$$=$$\frac{N(N+1)(2N+1)}{6}$ and $T_{s}f_{\Delta}$$=$$1$. The Fishier information matrix is
$\begin{split}{\left[\text{FIM}(\Theta)\right]}=4\pi^{2}\text{SNR}_{r}NL\\ \begin{bmatrix}\frac{1}{3}\frac{f_{\Delta}^{2}}{c^{2}}(N+1)(2N+1)&-\frac{1}{c\lambda_{c}}(L+1)(N+1)\\ -\frac{1}{c\lambda_{c}}(L+1)(N+1)&\frac{4}{3}\frac{T_{s}^{2}}{\lambda_{c}^{2}}(L+1)(2L+1)\end{bmatrix}.\end{split}$ (16)
Finally, the $\text{CRB}(\Theta)$ can be derived as
$\begin{split}{\text{CRB}(\Theta)}={\text{FIM}^{-1}(\Theta)}=\frac{1}{4\pi^{2}\text{SNR}_{r}NL\det({\text{FIM}(\Theta))}}\\ \begin{bmatrix}\frac{4}{3}\frac{T_{s}^{2}}{\lambda_{c}^{2}}(L+1)(2L+1)&\frac{1}{c\lambda_{c}}(L+1)(N+1)\\ \frac{1}{c\lambda_{c}}(L+1)(N+1)&\frac{1}{3}\frac{f_{\Delta}^{2}}{c^{2}}(N+1)(2N+1)\end{bmatrix}\\ =\frac{1}{\pi^{2}\text{SNR}_{r}NL(7NL-N-L-5)}\\ \begin{bmatrix}3T_{s}^{2}c^{2}\frac{(2L+1)}{(N+1)}&\frac{9}{4}\lambda_{c}c\\ \frac{9}{4}\lambda_{c}c&\frac{3}{4}\lambda_{c}^{2}f_{\Delta}^{2}\frac{(2N+1)}{(L+1)}\end{bmatrix}.\end{split}$ (17)

### v3

@@ III-C Target detection
The detection decision is based on the threshold, $\gamma$. The detector is given as $p\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}\gamma$, where, the detector indicates the presence of a target when $p>\gamma$, $\mathcal{H}_{1}$, and the absence of a target if $p<\gamma$, $\mathcal{H}_{0}$. Here, $p$ is defined as $p=\max_{r}\text{PAF}$.
@@ IV Cramér-Rao Bound
For simplicity, we consider the CRB of the random variables, range and velocity, denoted as $\Theta$$=$$[\Theta_{i},\Theta_{j}]^{T}$$=$$[d,v]^{T}$, in a single target case ($K$$=$$1$). We evaluate this for a given signal-to-noise ratio at BsB, $\text{SNR}_{r}={\rho\beta g^{2}}/({\sigma_{n}^{2}M})$, assuming a given RCS and fixed flight path. The received signal at the $n$th row, $l$th column of (7), disregarding the subscript $r,1$ and $k$, can be rewritten as
@@
where $g$$=$$\mathbf{a}^{T}(\theta_{d},\phi_{d})\mathbf{a}^{*}(\theta_{r},\phi_{r})$ is the beamforming gain of the $r$th transmitting beam with respect to the target’s angle. We consider the case with $M$=$100$ and $Q$=$81$ in (1) and (2). The gain $\left\lvert g\right\rvert$ is then bounded by $M(\text{dB})-3.7\text{dB}<\left\lvert g\right\rvert\leq M(\text{dB}),$ where $-3.7$ dB is from the selected $M$, and the bound holds irrespective of whether the target is located inside the sweeping beams, given that the $r$th beam is directed closest to the target among all $R$ beams. Formulate a vector $\mathbf{y}$$=$$[y_{1,1},y_{2,1},\dots,y_{2,1},y_{2,2},\dots,y_{N,L}]^{T},\;\mathbf{y}\in\mathbb{C}^{NL}$, from (9), i.e., by flattening $\mathbf{Y}$, where $[\mathbf{Y}]_{n,l}$$=$$y_{n,l}$. The Fisher information, $\text{FIM}(\Theta)$, for a complex Gaussian signal, as given in [12, Eq. (15.52)], is defined as
@@
where $\mathbf{C_{y}}$$=$$\sigma_{n}^{2}\mathbf{I}$ is the covariance matrix of $\mathbf{y}$. $\mathbf{\mu_{y}}$$=$$\sqrt{\frac{\rho}{M}\beta}\alpha g[e^{-j2\pi f_{\Delta}\frac{d}{c}}e^{j4\pi\frac{v}{\lambda_{c}}T_{s}},\dots,e^{-j2\pi Nf_{\Delta}\frac{d}{c}}e^{j4\pi\frac{v}{\lambda_{c}}LT_{s}}]^{T}$ is the mean of $\mathbf{y}$. Note that $\alpha$ is constant within an OFDM block length $L$ (Swerling 1 model).
@@ V Simulation Results
In Fig. 3, a target is generated with a random range, $d\leq d_{u}$, and random velocity, $\left\lvert v\right\rvert\leq v_{u}$. Fig. 3a shows CRB and range root mean square error (RMSE) of an $N^{\prime}$-point IFFT at SNRr$=$$-10$ dB, where the SSB line represents the received signal nullified at some symbols in the block, as shown in Fig. 1(b). Similarly, Fig. 3b shows the CRB and velocity RMSE of $L^{\prime}$-point FFT at SNRr$=$$-7$ dB. The results show that RMSE values for $N$$=$$240$, $L$$=$$4$, and the SSB signals are approximately the same for both cases, range and velocity. The absence of some symbols in the SSB signals has minimal effect compared to the completed block.
@@
Furthermore, the detection performance is investigated by uniformly and randomly deactivating different percentages of surveillance beams. To that, let us define $P_{fa}$$=$$P(\max_{r}\text{PAF}>\gamma|\mathcal{H}_{0})$ as the probability of false alarm and $P_{d}$$=$$P(\max_{r}\text{PAF}>\gamma|\mathcal{H}_{1})$ as the probability of detection. Thus, a $100$% beam deactivation indicates that all surveillance beams are turned off. Fig. 4 shows that with an appropriate thresholds i.e., $\gamma$$=$$4$, $P_{d}$ is maintained above $0.7$ up to $20$% beam deactivation with $P_{fa}$$=$$0.03$.
Figure 4: $P_{d}$ during beam deactivation (%).
@@ VI Conclusions
This paper investigates the feasibility of sensing low-altitude drones by employing the sweeping beams of SSB signals of the mobile network. A set of SSB signals is used as illuminators in a bistatic sensing setup (surveillance beams), and a base station serves as the sensing node. For detection, the decision relies on the threshold and is evaluated with the highest peak-to-average factor (PAF) of the 2D-FFT across the sweeping beams, while localization relies on the aggregated 2D-FFT, assisted by the PAF. Simulation results for the proposed scenario demonstrate the potential for drone-like target detection and localization.
@@ Appendix A Proof of proposition 1
where $[\text{FIM}(\Theta)]_{v,d}$$=$$[\text{FIM}(\Theta)]_{d,v}$, $\mathbb{E}\{\left\lvert\alpha\right\rvert^{2}\}$$=$$1$, $\text{SNR}_{r}$$=$${\rho\beta g^{2}}/({\sigma_{n}^{2}M})$, $\sum_{n=1}^{N}n$$=$${N(N+1)}/{2}$ ,$\sum_{n=1}^{N}n^{2}$$=$${N(N+1)(2N+1)}/{6}$ and $T_{s}f_{\Delta}$$=$$1$. The Fishier information matrix is
$\begin{split}{\text{FIM}(\Theta)}=4\pi^{2}\text{SNR}_{r}NL\\ \begin{bmatrix}\frac{1}{3}\frac{f_{\Delta}^{2}}{c^{2}}(N+1)(2N+1)&-\frac{1}{c\lambda_{c}}(L+1)(N+1)\\ -\frac{1}{c\lambda_{c}}(L+1)(N+1)&\frac{4}{3}\frac{T_{s}^{2}}{\lambda_{c}^{2}}(L+1)(2L+1)\end{bmatrix}.\end{split}$ (16)
Finally, the $\text{CRB}(\Theta)$ is derived as
$\begin{split}{\text{CRB}(\Theta)}={\text{FIM}^{-1}(\Theta)}=\frac{3}{\pi^{2}\text{SNR}_{r}NL(7NL-N-L-5)}\\ \begin{bmatrix}T_{s}^{2}c^{2}\frac{(2L+1)}{(N+1)}&\frac{3}{4}\lambda_{c}c\\ \frac{3}{4}\lambda_{c}c&\frac{1}{4}\lambda_{c}^{2}f_{\Delta}^{2}\frac{(2N+1)}{(L+1)}\end{bmatrix}.\end{split}$ (17)